# 1st Mehrdad Moradi H. Milton Stewart School of Industrial and Systems Engineering

### v1

@@ Abstract
Generative models have demonstrated significant success in anomaly detection and segmentation over the past decade. Recently, diffusion models have emerged as a powerful alternative, outperforming previous approaches such as GANs and VAEs. In typical diffusion-based anomaly detection, a model is trained on normal data, and during inference, anomalous images are perturbed to a predefined intermediate step in the forward diffusion process. The corresponding normal image is then reconstructed through iterative reverse sampling.
@@
We introduce Reconstruction-free Anomaly Detection with Attention-based diffusion models in Real-time (RADAR), which overcomes the limitations of reconstruction-based anomaly detection. Unlike current SOTA methods that reconstruct the input image, RADAR directly produces anomaly maps from the diffusion model, improving both detection accuracy and computational efficiency. We evaluate RADAR on real-world 3D-printed material and the MVTec-AD dataset. Our approach surpasses state-of-the-art diffusion-based and statistical machine learning models across all key metrics, including accuracy, precision, recall, and F1 score. Specifically, RADAR improves F1 score by 7% on MVTec-AD and 13% on the 3D-printed material dataset compared to the next best model.
Code available at: https://github.com/mehrdadmoradi124/RADAR
Index Terms:
Unsupervised Anomaly Detection, Diffusion Models, Additive Manufacturing
@@ I Introduction
With significant advancements in sensing technology, the volume of image data collected for quality control has increased exponentially over the past decade. This rapid growth has created a strong demand for reliable, real-time algorithms capable of performing online anomaly detection and segmentation. Such capabilities are crucial across diverse industries, including stone countertop manufacturing [1], ceramic capacitor chip production [2], softwood lumber grading [3], material microstructure monitoring [4, 5], fabric inspection [6], and three-dimensional printing for additive manufacturing [7].
@@
The structure of this paper is as follows. We review the related literature on image-based anomaly detection for patterned data in Section II, present the problem formulation and our proposed methodology in Section III, evaluate the effectiveness of RADAR through extensive experiments in Section IV, and conclude the paper in Section V. Our key contributions are summarized below:
@@ II-A Statistical and Deep Learning Methods
Convolutional neural networks (CNNs) have been widely used for anomaly detection and segmentation [14, 15, 16]. More recently, researchers have explored the use of pre-trained models to extract and compare image embeddings. For example, [23, 24] applied Gaussian mixture models to the embeddings derived from pre-trained networks and detected anomalies by measuring the Mahalanobis distance between the test image embedding and the distribution of normal embeddings. Similarly, [25] trained a neural network to learn mappings between embeddings from two different pre-trained models and identified anomalies using the mean squared error between predicted and actual embeddings.
@@ II-B Diffusion Models for Anomaly Detection
Diffusion models have demonstrated impressive performance in image synthesis [20]. In the context of anomaly detection, a diffusion model is trained exclusively on normal data and then used to reconstruct anomalous samples by generating their corresponding “normal” counterparts. These models can be viewed as approximating a stochastic differential equation (SDE) in both the forward and reverse diffusion processes [26], making the sampling path inherently stochastic. Guidance during the reverse sampling process is often necessary to ensure that the reconstructed image preserves the structural features of the input anomaly image. A key limitation of diffusion models is their high computational cost, as achieving high-quality reconstructions typically requires hundreds of reverse steps [17]. To address this, latent diffusion models [27] operate in the lower-dimensional embedding space of an encoder, reducing computational complexity while maintaining fidelity.
@@ III-A Problem Formulation and Overview
Given a set of normal training images $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, where N is the total number of training samples, our goal is to learn a model f that can determine whether a new test image $\hat{x}$ is anomalous (label =1) or normal (label =0) based on its deviation from the distribution of normal data. Importantly, this is achieved without access to pixel-level anomaly annotations, which are expensive to collect, or any anomalous images during training.
To construct f, we train a diffusion U-Net $\epsilon_{\theta}$ exclusively on normal data. During inference, we compute a compact feature representation by applying a Sobel edge detector and the L2 norm to the outputs of $\epsilon_{\theta}$ on noisy versions of the input image. The Sobel operator highlights local structural changes (e.g., edges), while the L2 norm aggregates these variations into a feature vector. An overview of our method is shown in Figure 1, and each step is detailed in the following subsections.
@@ III-B Diffusion Model Background
Denoising Diffusion Probabilistic Models (DDPMs) [17] are a class of generative models that learn data distributions by gradually adding Gaussian noise in a forward process and then reversing this process. The forward diffusion process is defined as a Markov chain that incrementally perturbs the data:
@@
where $t\in\{1,\ldots,T\}$ is the time step, $\mathbf{x}_{t}$ is the noisy sample at step t, and $\beta_{t}$ is the predefined noise schedule. Figure 2 provides an overview of this forward and backward process.
@@
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. This loss effectively trains $\epsilon\theta$ to predict the noise added at each step, commonly implemented with a U-Net architecture [37], as illustrated in Figure 3.
@@ III-C Single-step Anomaly Map Generation
When using diffusion models for reconstruction, the anomalous image is first corrupted by adding noise up to a certain level. This noise injection serves two purposes: (1) it disrupts anomaly pixels so they do not influence the reverse generation process, and (2) it preserves normal pixels sufficiently to act as implicit guidance during backward diffusion. However, achieving this balance is challenging—especially when anomaly pixels exhibit smaller intensity values than normal patterns—because no single noise step can both erase anomalies and retain fine normal features. Furthermore, reconstruction-based approaches are computationally expensive due to the need for multiple iterative sampling steps in reverse diffusion sometimes 500 steps. In low-data engineering applications, they often overfit and fail to accurately reconstruct complex patterns.
@@
An example output of the predicted noise is shown in Figure 1. Due to the distribution shift, the predicted noise behaves very differently in anomalous pixels but remains consistent in normal regions, regardless of the underlying normal pattern. In the next step, features are extracted from this noise map.
@@ III-D Feature Extraction
Figure 1 visualizes the two-dimensional feature space for both in-control and out-of-control data from Phase 2 of our first case study on 3D printed material images. The features reveal a clear separation between the two classes, paving the way for effective binary classification in the next step.
@@ III-F Pixel Level Anomaly Map
The single-step nature of our algorithm enables real-time implementation across various engineering applications. Once a sample is identified as anomalous, pixel-level localization is achieved by applying Sobel edge detection to a blurred version of the image, where blurring reduces noise and emphasizes anomaly boundaries. The Sobel output highlights the pixels corresponding to anomalous regions. Representative qualitative examples are shown in Figures 6 and 7.
@@ IV-A1 Datasets
We use the dataset introduced in [7], which contains 3D printed material images from the Fused Filament Fabrication (FFF) process with two distinct deposition strategies (45° and 135°). These deposition angles are chosen to represent different structural patterns in the printed material. Examples of defective samples are shown in the left column of Figure 6. Additionally, we use 107 images from the tile category of the MVTec-AD dataset [43], which contains highly stochastic surface patterns. These textures complement the periodic patterns observed in the 3D Print dataset. The specifications of these two datasets are summarized in Table I.
Dataset Category Count
3D Prints Training (Normal): 135° / 45° 40 / 41
Total Training Images 81
Total Training Patches 431,649
Test (Normal): 135° / 45° 19 / 21
Test (Anomaly): 135° / 45° 22 / 22
Total Test Images 84
Tile Training (Normal) 230
Total Training Patches 1,225,670
Test (Normal) 33
Test (Anomaly): Crack / Glue 15 / 16
Test (Anomaly): Gray / Oil / Rough 14 / 16 / 13
Total Test Images 107
TABLE I: Compact summary of dataset specifications for 3D Prints and Tile datasets, including total number of training patches.
@@ IV-A2 Implementation Details
The U-Net architecture we use closely follows the design of the stable diffusion model [27]. As shown in Figure 5, it consists of a downsampling block, an upsampling block, and a mid block, all connected through residual and skip connections. Each block is composed of ResNet blocks, self-attention modules, and either downsampling or upsampling layers. For downsampling, a 2D convolution with a kernel size of 4, stride of 2, and padding of 1 is applied to halve the spatial resolution. Conversely, for upsampling, we use a transposed 2D convolution (ConvTranspose2d) with the same kernel size, stride, and padding to double the resolution. Each ResNet block includes a normalization layer, SiLU activation functions, and convolutional operations. The self-attention module leverages learnable keys, queries, and values to compute attention weights across all spatial positions, enabling the model to capture long-range dependencies. This mechanism allows the U-Net to focus on the most informative regions of the feature maps, improving its ability to detect subtle anomalies that may not be captured by purely convolutional operations.
For the diffusion model, we adopt the Denoising Diffusion Probabilistic Model (DDPM) with 1000 forward diffusion steps, following the setup in the original work [17]. The noise schedule linearly increases from 0.0001 to 0.02. We train the model using a batch size of 64 and a learning rate of 0.0001. All models were trained using an NVIDIA RTX6000 GPU.
Figure 5: U-Net architecture. The network consists of a downsampling block, a middle block, and an upsampling block. Each block integrates attention modules, convolutional layers, and residual connections, which are represented by arrows with a plus sign.
@@ IV-A4 Evaluation Metrics
We evaluate the performance of our model using accuracy, precision, recall, and F1 score. Accuracy measures the overall correctness of the model’s predictions but does not differentiate between the performance on positive and negative classes. Precision evaluates the proportion of correctly identified anomalies among all predicted anomalies. Recall, on the other hand, measures the proportion of actual anomalies correctly detected by the model. The F1 score, defined as the harmonic mean of precision and recall, provides a balanced metric that captures both false positives and false negatives.
IV-B Anomaly Detection Comparison
Table II presents the performance of RADAR compared to competing methods. On the 3D printed material dataset, RADAR consistently outperforms all other models across all metrics. Notably, reconstruction-based diffusion models perform poorly, highlighting their limitations when anomaly patterns closely resemble normal patterns in pixel intensity. The best-performing diffusion-based method, DiffusionAD, achieves an F1 score of only 0.27—well below the 0.5 F1 score expected from a random classifier in a balanced binary setting. In fact, AnoDDPM and DiffusionAD, despite incorporating simplex noise and norm-guided reconstruction respectively, underperform compared to the baseline DDPM. In contrast, the statistical machine learning models L&F and B&A show strong performance, with B&A matching RADAR’s precision (0.77) and achieving the second-highest F1 score of 0.72, trailing RADAR’s 0.85.
Compared to the 3D Print dataset, the metrics on the Tile dataset are overall lower, indicating that its stochastic patterns pose a greater challenge for our method. Interestingly, unlike with 3D Prints, reconstruction-based diffusion models outperform the statistical approaches on this dataset. AnoDDPM achieves the second-best F1 score of 0.60 (compared to RADAR’s 0.67), suggesting that the simplex noise mechanism in AnoDDPM is particularly effective for stochastic patterns. As illustrated in Figure 7, anomalies in the Tile dataset exhibit clear pixel-intensity differences from normal patterns, which likely benefits diffusion-based methods. However, DiffusionAD performs poorly, implying that its norm-guided, two-step reconstruction is unsuitable for complex stochastic textures.Another notable observation is the poor performance of L&F compared to B&A, despite their conceptual similarities. This difference may stem from how the anomaly statistics are computed in L&F. For this dataset, RADAR achieves the highest F1 score and accuracy overall. While L&F attains the highest precision (0.95), its recall is only 0.07, indicating zero false positives but many missed anomalies in the from of false negatives. Conversely, DDPM and AnoDDPM achieve the highest recall (0.57), slightly outperforming RADAR’s recall of 0.51.
@@ IV-B Anomaly Detection Comparison
TABLE II: Comparison of methods across Accuracy, Precision, Recall, and F1 Score for 3D Prints and Tile datasets.
@@ IV-C Pixel Level Anomaly Segmentation
The qualitative results shown in Figures 6 and 7 present the segmentation masks of sample images identified as anomalous by the algorithm. In Figure 6, both 45° and 135° deposition patterns with various types of anomalies are visualized, where the anomaly heatmaps on the right accurately localize the anomalous pixels. Figure 7 illustrates anomalies such as cracks and gray strokes on tile images, along with their corresponding heatmaps.
@@ IV-D Sensitivity Analysis
As discussed earlier, the contamination level must be set when using the Isolation Forest algorithm for one-class classification. In our experiments, we fixed this parameter at 0.05. To evaluate its impact, we varied the contamination level from 0 to 0.5 and plotted all metrics for each diffusion-based model on the Tile dataset. As illustrated in Figure 8, our model consistently outperforms all other methods in terms of accuracy, precision, and F1 score, though it slightly underperforms in recall compared to DDPM and AnoDDPM.

### v2

# Single-Step Reconstruction-Free Anomaly Detection and Segmentation via Diffusion Models

@@ Abstract
Generative models have demonstrated significant success in anomaly detection and segmentation over the past decade. Recently, diffusion models have emerged as a strong alternative, outperforming previous approaches such as GANs and VAEs. In typical diffusion-based anomaly detection, a model is trained on normal data, and during inference, anomalous images are perturbed to a predefined intermediate step in the forward diffusion process. The corresponding normal image is then reconstructed through iterative reverse sampling.
@@
We introduce Reconstruction-free Anomaly Detection with Attention-based diffusion models in Real-time (RADAR), which overcomes the limitations of reconstruction-based anomaly detection. Unlike current state-of-the-art (SOTA) methods that reconstruct the input image, RADAR directly produces anomaly maps from the diffusion model, improving both detection accuracy and computational efficiency. We evaluate RADAR on real-world 3D-printed material and the MVTec-AD dataset. Our approach surpasses state-of-the-art diffusion-based and statistical machine learning models across all key metrics, including accuracy, precision, recall, and F1 score. Specifically, RADAR improves F1 score by 7% on MVTec-AD and 13% on the 3D-printed material dataset compared to the next best model.
Code: https://github.com/mehrdadmoradi124/RADAR
@@ I Introduction
With significant advancements in sensing technology, the volume of image data collected for quality control has increased exponentially over the past decade. This rapid growth has created an increased demand for reliable, real-time algorithms capable of performing online anomaly detection and segmentation. Such capabilities are crucial across diverse industries, including stone countertop manufacturing[19], ceramic capacitor chip production[17], softwood lumber grading[3], material microstructure monitoring[20, 33], fabric inspection[5], and three-dimensional printing for additive manufacturing[6].
@@
The structure of this paper is as follows. We review the related literature on image-based anomaly detection for patterned data in Sec. II, present the problem formulation and our proposed methodology in Sec. III, evaluate the effectiveness of RADAR through extensive experiments in Sec. IV, and conclude the paper in Sec. V. Our key contributions are summarized below:
@@ II-A Statistical and Deep Learning Methods
Convolutional neural networks (CNNs) have been frequently used for anomaly detection and segmentation[37, 24, 14]. More recently, researchers have explored the use of pre-trained models to extract and compare image embeddings. For example, [34, 21] applied Gaussian mixture models to the embeddings derived from pre-trained networks and detected anomalies by measuring the Mahalanobis distance between the test image embedding and the distribution of normal embeddings. Similarly, [35] trained a neural network to learn mappings between embeddings from two different pre-trained models and identified anomalies using the mean squared error between predicted and actual embeddings.
@@ II-B Diffusion Models for Anomaly Detection
Diffusion models have demonstrated strong performance in image synthesis[8]. In the context of anomaly detection, a diffusion model is trained exclusively on normal data and then used to reconstruct anomalous samples by generating their corresponding “normal” counterparts. If the training data are contaminated, RDDPM[22] uses the Huber loss to achieve robustness to outliers. These models can be viewed as approximating a stochastic differential equation (SDE) in both the forward and reverse diffusion processes[31], making the sampling path inherently stochastic. Guidance during the reverse sampling process is often necessary to ensure that the reconstructed image preserves the structural features of the input anomaly image. A key limitation of diffusion models is their high computational cost, as achieving high-quality reconstructions typically requires hundreds of reverse steps[12]. To address this, latent diffusion models[28] operate in the lower-dimensional embedding space of an encoder, reducing computational complexity while maintaining fidelity.
@@ III-A Problem Formulation and Overview
Given a set of normal training images $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, where N is the total number of training samples, our objective is to learn a model f that can determine whether a new test image $\hat{x}$ is anomalous (label =1) or normal (label =0) based on its deviation from the distribution of normal data. Importantly, this is achieved without access to pixel-level anomaly annotations, which are expensive to collect, or any anomalous images during training.
To construct f, we train a diffusion U-Net $\epsilon_{\theta}$ exclusively on normal data. During inference, we compute a compact feature representation by applying a Sobel edge detector and the L2 norm to the outputs of $\epsilon_{\theta}$ on noisy versions of the input image. The Sobel operator highlights local structural changes (e.g., edges), while the L2 norm aggregates these variations into a feature vector. An overview of our method is shown in Fig. 1, and each step is detailed in the following subsections.
@@ III-B Diffusion Model Background
Denoising Diffusion Probabilistic Models (DDPMs)[12] are a class of generative models that learn data distributions by gradually adding Gaussian noise in a forward process and then reversing this process. The forward diffusion process is modeled as a Markov chain that incrementally perturbs the data:
@@
where $t\in\{1,\ldots,T\}$ is the time step, $\mathbf{x}_{t}$ is the noisy sample at step t, and $\beta_{t}$ is the predefined noise schedule. Fig. 2 provides an overview of this forward and backward process.
@@
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. This loss effectively trains $\epsilon\theta$ to predict the noise added at each step, commonly implemented with a U-Net architecture[29], as illustrated in Fig. 3.
@@ III-C Single-step Anomaly Map Generation
When using diffusion models for reconstruction, the anomalous image is first corrupted by adding noise up to a certain level. This noise injection serves two purposes: (1) it disrupts anomaly pixels so they do not influence the reverse generation process, and (2) it preserves normal pixels sufficiently to act as implicit guidance during backward diffusion. However, achieving this balance is challenging—especially when anomaly pixels exhibit smaller intensity values than normal patterns—because no single noise step can both erase anomalies and retain fine normal features. Furthermore, reconstruction-based approaches are computationally expensive due to the need for multiple iterative sampling steps in reverse diffusion, sometimes up to 500 steps. In low-data engineering applications, they often overfit and fail to accurately reconstruct complex patterns.
@@
An example output of the predicted noise is shown in Fig. 1. Due to the distribution shift, the predicted noise behaves very differently in anomalous pixels but remains consistent in normal regions, regardless of the underlying normal pattern. In the next step, features are extracted from this noise map.
@@ III-D Feature Extraction
Fig. 1 visualizes the two-dimensional feature space for both in-control and out-of-control data from Phase 2 of our first case study on 3D printed material images. The features reveal a clear separation between the two classes, paving the way for effective binary classification in the next step.
@@ III-F Pixel Level Anomaly Map
The single-step nature of our algorithm enables real-time implementation across various engineering applications. Once a sample is identified as anomalous, pixel-level localization is achieved by applying Sobel edge detection to a blurred version of the image, where blurring reduces noise and emphasizes anomaly boundaries. The Sobel output highlights the pixels corresponding to anomalous regions. Representative qualitative examples are shown in Figs. 6 and 7.
@@ IV-A1 Datasets
We use the dataset introduced in [6], which contains 3D printed material images from the Fused Filament Fabrication (FFF) process with two distinct deposition strategies (45° and 135°). These deposition angles are chosen to represent different structural patterns in the printed material. Examples of defective samples are shown in the left column of Fig. 6. Additionally, we use 107 images from the tile category of the MVTec-AD dataset[2], which contains highly stochastic surface patterns. These textures complement the periodic patterns observed in the 3D Print dataset.
3D Prints (FFF) contains two deposition patterns, $45^{\circ}$ and $135^{\circ}$: training has 81 normal images (40 @ $135^{\circ}$, 41 @ $45^{\circ}$; testing has 84 images—40 normal (19 @ $135^{\circ}$, 21 @ $45^{\circ}$) and 44 anomalous (22 @ $135^{\circ}$, 22 @ $45^{\circ}$). MVTec-AD Tile has 230 normal training images and 107 test images comprising 33 normal and 74 anomalous across five defect types: crack (15), glue strip (16), gray stroke (14), oil (16), and rough (13).
@@ IV-A2 Implementation Details
The U-Net architecture we use closely follows the design of the stable diffusion model[28]. As shown in Fig. 5, it consists of a downsampling block, an upsampling block, and a mid block, all connected through residual and skip connections. Each block is composed of ResNet blocks, self-attention modules, and either downsampling or upsampling layers. For downsampling, a 2D convolution with a kernel size of 4, stride of 2, and padding of 1 is applied to halve the spatial resolution. Conversely, for upsampling, we use a transposed 2D convolution (ConvTranspose2d) with the same kernel size, stride, and padding to double the resolution. Each ResNet block includes a normalization layer, SiLU activation functions, and convolutional operations. The self-attention module leverages learnable keys, queries, and values to compute attention weights across all spatial positions, enabling the model to capture long-range dependencies. This mechanism allows the U-Net to focus on the most informative regions of the feature maps, improving its ability to detect subtle anomalies that may not be captured by purely convolutional operations.
For the diffusion model, we adopt the Denoising Diffusion Probabilistic Model (DDPM) with 1000 forward diffusion steps, following the setup in the original work[12]. The noise schedule linearly increases from 0.0001 to 0.02. Training is done with a batch size of 64 and a learning rate of 0.0001. All models were trained using an NVIDIA RTX6000 GPU.
Figure 5: U-Net architecture. The network consists of a downsampling block, a middle block, and an upsampling block. Each block integrates attention modules, convolutional layers, and residual connections, which are represented by arrows with a plus sign. Convolution, transposed convolution and multi-head attention pictures were generated by GPT-4o [25].
@@ IV-A4 Evaluation Metrics
The performance of our model is evaluated with accuracy, precision, recall, and F1 score. Accuracy shows the overall correctness of the model’s predictions but does not differentiate between the performance on positive and negative classes. Precision measures the proportion of correctly detected anomalies among all predicted anomalies. Recall, on the contrary, measures the proportion of actual anomalies correctly detected by the model. As the harmonic mean of precision and recall, the F1 score provides a balanced metric that reflects both false positives and false negatives.
IV-B Anomaly Detection Results
Tab. I presents the performance of RADAR compared to competing methods. On the 3D printed material dataset, RADAR consistently outperforms all other models across all metrics. Notably, reconstruction-based diffusion models perform poorly, highlighting their limitations when anomaly patterns closely resemble normal patterns in pixel intensity. The best-performing diffusion-based method, DiffusionAD, achieves an F1 score of only 0.27—well below the 0.5 F1 score expected from a random classifier in a balanced binary setting. In fact, AnoDDPM and DiffusionAD, despite incorporating simplex noise and norm-guided reconstruction respectively, underperform compared to the baseline DDPM. In contrast, the statistical machine learning models L&F and B&A show strong performance, with B&A matching RADAR’s precision (0.77) and achieving the second-highest F1 score of 0.72, trailing RADAR’s 0.85.
Compared to the 3D Print dataset, the metrics on the Tile dataset are overall lower, indicating that its stochastic patterns pose a greater challenge for our method. Interestingly, unlike with 3D Prints, reconstruction-based diffusion models outperform the statistical approaches on this dataset. AnoDDPM achieves the second-best F1 score of 0.60 (compared to RADAR’s 0.67), suggesting that the simplex noise mechanism in AnoDDPM is particularly effective for stochastic patterns. As illustrated in Fig. 7, anomalies in the Tile dataset exhibit clear pixel-intensity differences from normal patterns, which likely benefits diffusion-based methods. However, DiffusionAD performs poorly, implying that its norm-guided, two-step reconstruction is unsuitable for complex stochastic textures. Another notable observation is the poor performance of L&F compared to B&A, despite their conceptual similarities. This difference may stem from how the anomaly statistics are computed in L&F. For this dataset, RADAR achieves the highest F1 score and accuracy overall. While L&F attains the highest precision (0.95), its recall is only 0.07, indicating zero false positives but many missed anomalies in the form of false negatives. Conversely, DDPM and AnoDDPM achieve the highest recall (0.57), slightly outperforming RADAR’s recall of 0.51.
@@ IV-B Anomaly Detection Results
TABLE I: Comparison of methods across Accuracy, Precision, Recall, and F1 Score for 3D Prints and Tile datasets.
@@ IV-C Pixel Level Anomaly Segmentation
The qualitative results shown in Figs. 6 and 7 present the segmentation masks of sample images identified as anomalous by the algorithm. In Fig. 6, both 45° and 135° deposition patterns with various types of anomalies are visualized, where the anomaly heatmaps on the right accurately localize the anomalous pixels. Fig. 7 illustrates anomalies such as cracks and gray strokes on tile images, along with their corresponding heatmaps.
@@ IV-D Sensitivity Analysis
As discussed earlier, the contamination level must be set when using the Isolation Forest algorithm for one-class classification. In our experiments, we fixed this parameter at 0.05. To evaluate its impact, we varied the contamination level from 0 to 0.5 and plotted all metrics for each diffusion-based model on the Tile dataset. As illustrated in Fig. 8, our model consistently outperforms all other methods in terms of accuracy, precision, and F1 score, though it slightly underperforms in recall compared to DDPM and AnoDDPM.